A certain radioactive element decays to 1/2 of its original amount every 20 years. If there was 100mg to start with, how much will be present after 200 years? Set up an equation and solve the problem. Can someone please help me ??
Answer by stanbon(75887)

A certain radioactive element decays to 1/2 of its original amount every 20 years. If there was 100mg to start with, how much will be present after 200 years? Set up an equation and solve the problem.
------------------
A(t)=A(0)*(1/2)^(t/20)
A(t) is the amount you have at a certain time.
A(0) is the amount you have when you start.
---------------------------
A(200)=100*(1/2)^(200/20)
A(200) = 100*0.0313
A(200) = 31.3 mg
===================
Cheers,
Stan H.